OpenStudy (anonymous):

I need help!! rewrite using only positive rational exponents (x^6/7)^-7/5 simplify your answer.

OpenStudy (anonymous):

Use the rules: \[ x^{-a}=\frac{1}{x^a}\\ (x^a)^b=x^{ab} \]

OpenStudy (anonymous):

is it x^6/5

OpenStudy (anonymous):

It'd be: \[ \left(x^\frac{6}{7}\right)^{-\frac{7}{5}}=x^{-\frac{6}{5}}=\frac{1}{x^\frac{6}{5}} \]